Convert y = 2 (16/9)^3x to base 16/9
1
Expert's answer
2021-04-21T18:05:35-0400

Given, the function "y=2(16\/9)^{3x}" .

Divide by 2 both side.

"y=2(16/9)^{3x}
\newline
\frac{y}{2}=(\frac{16}{9})^{3x}"

Apply logarithm both side.

"log(\frac{y}{2})=3xlog(\frac{16}{9})"

Divide by "log(\frac{16}{9})" both side.

"\frac{log(\frac{y}{2})}{log(\frac{16}{9})}=3x"

From logarithm property, "log_{b} a=\frac{log_{e}a}{log_{e}b}".

Then, "log_{\frac{16}{9}}\frac{y}{2}=3x".

Thus, the required answer is "log_{\frac{16}{9}}\frac{y}{2}=3x".